(a) Find the area enclosed by the curve 𝑦 = 25 − 𝑥2

and the straight line 

𝑦 = 𝑥 + 13 correct to 1 decimal place.

(b)Calculate the arc length of the curve 𝑦 = 𝑥

3

2 between 𝑥 = 0 𝑎𝑛𝑑 𝑥 = 1

correct to 1 decimal place


1
Expert's answer
2021-03-24T07:41:40-0400

To get the points of intersection,

y=y

"25-x^2=x+13\\
x^2+x-12=0\\
x=-4, x=3"

The 2 points of intersection are at x = -4 and x = 3.

"y=25- x^2\\
INT(x) = 25x - {\frac{x^{3}}{3}}\\
INT(3) = 75 - 9 = 66\\
INT(x) = 25x - {\frac{x^{3}}{3}}\\
INT(-4) = -100 + \frac{64}{3} = \frac{-236}{3}"

Area under parabola to x-axis,

"= 66 +\frac{236}{3} = \frac{434}{3}\\

\text{Area under line}= 87.5 sq unit.\\
b)\\
F(x)=x^{\frac{3}{2}}\\
(f'(x))^2=9x\\
\text{Arc length}=\int_0^1 \sqrt{1+9x}dx\\
u=1+9x, du=9dx\\
When, x=0, u=1, When, x=1, u=10\\
={\frac{1}{9}}\int_0^1\sqrt{1+9x}dx\\
={\frac{1}{9}}\int_1^{10}\sqrt{u}du\\
={\frac{1}{9}}\times{\frac{2}{3}} u^{\frac{3}{2}}|_1^{10}\\
=2.3units"